# A review of regularised estimation methods and cross-validation in spatiotemporal statistics

### v1

@@ Abstract
This review article focuses on regularised estimation procedures applicable to geostatistical and spatial econometric models. These methods are particularly relevant in the case of big geospatial data for dimensionality reduction or model selection. To structure the review, we initially consider the most general case of multivariate spatiotemporal processes (i.e., $g>1$ dimensions of the spatial domain, a one-dimensional temporal domain, and $q\geq 1$ random variables). Then, the idea of regularised/penalised estimation procedures and different choices of shrinkage targets are discussed. Finally, guided by the elements of a mixed-effects model, which allows for a variety of spatiotemporal models, we show different regularisation procedures and how they can be used for the analysis of geo-referenced data, e.g. for selection of relevant regressors, dimensionality reduction of the covariance matrices, detection of conditionally independent locations, or the estimation of a full spatial interaction matrix.
Keywords: Big data, geostatistics, regularised estimation, review, spatiotemporal statistics
@@ 1 Introduction
In the era of big geospatial data, the analysis of intricate spatial and spatiotemporal processes has become both increasingly vital and challenging. This is attributed to the growing volume and finer resolution of geo-referenced data, including remotely sensed and LiDAR data, as well as the increasing diversity of data types due to automated collection and processing, such as social network data and image data. These advancements enable the examination of more intricate relationships and allow for the detection of weaker dependencies or variations in the data. To statistically investigate such interactions, geostatistical and spatial econometric models offer a robust framework for comprehending the fundamental structures of these processes. Additionally, they provide insights into spatial dependencies, temporal dynamics, and the interactions among multiple random variables. An essential advantage of these statistical models lies in their interpretability, unlike deep learning models, which are often viewed as black-box models unless explicitly designed for interpretability. However, as the dimensionality of the spatial domain increases and datasets become larger and more complex, conventional modelling approaches frequently encounter challenges related to computational complexity and the appropriate selection of influential variables.
@@
Regularised estimation procedures offer effective solutions for these challenges from multiple perspectives. For instance, they can be used for model selection and dimensionality reduction but also to reveal spatial dependence structures going beyond geographical proximity. This article will provide a comprehensive review of regularisation techniques, including shrinkage and penalisation methods, used for statistical modelling of geospatial data. Thereby, our focus will be on geostatistical models (Section 3) and spatial econometric models (Section 4). Since the degree of regularisation is typically chosen based on predictive accuracy, we also review cross-validation techniques which are suitable under spatiotemporal dependence (Section 5). Finally, Section 6 concludes the article and discusses some direction for future research.
@@ 2 General framework
Geo-referenced $q$-variate data across space and time can be considered as a set of random variables $\{Y_{t}(\boldsymbol{s})\in\mathds{R}^{q}:\boldsymbol{s}\in D_{\boldsymbol{s}},t\in D_{t}\}$ that are observed with a certain order in time and space. The spatial domain is denoted by $D_{\boldsymbol{s}}\subseteq\mathds{R}^{g}$, which is a subspace with positive volume of the $g$-dimensional real space with $g\geq 2$ (see Cressie and Wikle,, 2015). By contrast, the time domain is one-dimensional, i.e., $D_{t}\subseteq\mathds{R}$, typically assumed to be a subset of integers $D_{t}\subseteq\mathds{Z}$, implying that the data are equidistantly observed in time. For $D_{t}$ being a singleton $\{T\}$, i.e., one single time point $T$, the process collapses to a purely spatial process without a temporal dimension. Moreover, if $D_{\boldsymbol{s}}$ is a singleton of one location only, we get a $q$-variate time series process. There are different sources of (statistical) dependence for such geo-referenced data:
@@
    Temporal proximity between two observations typically causes correlations. That is, the closer two observations are in the temporal domain $D_{t}$, the higher they are correlated, e.g., the current outside temperature will be similar to the temperature one hour ago but less similar to the temperature 12 hours ago.
@@
    Geographical proximity in $D_{\boldsymbol{s}}$ induces a similar dependence, known as spatial dependence. As claimed by Tobler’s first law of Geography, two observations are more similar if they are close to each other. That is, the current outside temperature in a city is similar to that in the neighbouring cities but less similar if you move away, especially in the North-South direction.
@@
There are two (competing) approaches in spatial and spatiotemporal statistics to account for spatial dependence. Firstly, the dependence can be modelled using spatially dependent processes, where a suitable covariance function defines the entries of the covariance matrix. This approach is commonly known as geostatistics. Secondly, the dependent/outcome variable can be explicitly correlated with the nearby observations. Thereby, the local neighbourhood is defined by a suitable weight matrix $\mathbf{W}$, which is weighting all other observations, such that the product of $\mathbf{W}$ and the dependent variable is the weighted average of the adjacent observations. This approach is usually called the spatial econometrics approach, even though it is not tied to economic applications. Thus, we prefer to refer to this approach as spatial autoregression. Below, we will provide an overview of regularised estimation procedures for both approaches, starting with geostatistical models in Section 3 and followed by spatial autoregression in Section 4.
@@ 3 Geostatistical models
Generally, there are different sources for increasing computational complexity. Firstly, if the number of spatial locations $n$ increases, the covariance matrices of the process increase quadratically in $n$. It is worth noting that an increasing length of the time series is typically less critical because there is a causal relation in time (i.e., only past observations can influence future observations), such that spatiotemporal processes can be modelled conditionally on the past realisations. Secondly, if the number of variables $q$ increases, we observe the same quadratic increase in the covariances between all variables. While the spatial correlation typically decreases with an increasing distance between the observations, which often allows for a sparse representation of the covariance matrices; the cross-correlation matrices are often dense.
Consider a mixed-effects spatiotemporal model, that is,
@@
where $\mu_{t}(\boldsymbol{s})$ is the fixed-effects model, $\omega_{t}(\boldsymbol{s})$ is the random-effects model, and $\varepsilon_{t}(\boldsymbol{s})$ is the modelling error at time point $t$ and location $\boldsymbol{s}$. While the fixed-effects term typically models the mean behaviour and the influence of exogenous variables, the random-effects term accounts for any additional dependence in the data, e.g., temporal and spatial dependence, heterogeneity, or cross-correlation between the variables. Now, let us distinguish between regularised estimation procedures for the parameters of $\mu_{t}(\boldsymbol{s})$ and $\omega_{t}(\boldsymbol{s})$.
Penalised estimation procedures can be used to shrink certain model parameters towards a pre-specified target. The most often applied shrinkage target is zero. In this case, the parameters are shrunk towards zero, which means that the parameter is excluded from the model if the estimated coefficient is equal to the zero target. Thus, these methods are suitable for model selection. When applied to the fixed effects model, it can be used for automatically selecting relevant regressors. In the following Section 3.1, we will focus on this in more detail. Moreover, a shrinkage target of zero can be used to introduce zeros in the spatial covariance matrix or the precision matrix, which would indicate the conditional independence between these observations. This will be the subject of Section 3.2.
3.1 Regularisation for the large-scale variation
@@
Linear regression models, whether referring to spatiotemporal or unstructured data, can be estimated by various techniques, including Likelihood maximisation and least squares. However, in high-dimensional settings, traditional methods for regression models might not be directly applied (Nandy et al.,, 2017). In order to reduce the model’s complexity due to the fixed effect component $\mu_{t}(\boldsymbol{s})$, several methods for spatial and spatiotemporal settings have been proposed. Particularly, we refer to the task of selecting the most relevant predictors among a large set of candidates.
@@
The state-of-the-art literature inherent to the statistical (algorithmic) variable selection task pivots around the Least Absolute Shrinkage and Selection Operator (LASSO) approach introduced by Tibshirani, (1996). LASSO is a penalised version of ordinary least squares that uses a LASSO-type $l_{1}$ penalisation to shrink irrelevant parameters to zero. Under mild regularity conditions, including uncorrelated observations, LASSO ensures consistent parameter estimation (see Section 2.4.2 of Bühlmann and Van De Geer,, 2011) and model selection consistency (Bickel et al.,, 2009; Belloni and Chernozhukov,, 2013), i.e., LASSO owns the oracle property. Among others, one of the most relevant features of LASSO is that it admits solutions to the minimisation task even when the number of parameters is greater than the number of available observations. Also, LASSO penalty (and extensions) can be used either considering least squares and penalised likelihood problems (Fan and Li,, 2001).
@@
It is worth noting that real-world geostatistical applications are prone to cross-correlated regressors due to their dependent (spatially and temporally) structure. Zhao and Yu, (2006) point out that when predictors are correlated, the classic LASSO algorithm does not provide selection-consistent estimates. Furthermore, when cross-correlation is detected, also group-LASSO estimators, which assume orthonormal data within each group, perform poorly in selecting the relevant predictors (Simon and Tibshirani,, 2012). A straightforward solution to this issue is provided by the adaptive LASSO penalty, which leads to selection-consistent estimators even in the presence of cross-correlated covariates (see Zou,, 2006) and Zou and Li, (2008). Among others, Wen et al., (2018) use a spatial autoregressive model with adaptive LASSO penalty to detect relevant autoregressive parameters in genetic studies. Zhu et al., (2010) implemented an iterative penalised likelihood estimator with adaptive LASSO penalty to select predictors and neighbourhood structure in conditional autoregressive and simultaneous autoregressive models with spatially correlated error terms. Later on, Reyes et al., 2012a proposed an adaptive LASSO algorithm for the case of linear regression models with spatiotemporal neighbourhood structures. Liu, (2017) extended the previous algorithms, allowing for spatial correlation to be captured by either the spatial lag terms or spatial errors or both through a SARAR model. Also, their penalised estimates are obtained via least squares approximation to account for possible non-concavity of the likelihood function. Other examples of penalised likelihood for spatiotemporal data are in Al-Sulami et al., (2019), in which an adaptive LASSO method is proposed to simultaneously identify and estimate spatiotemporal lag interactions in the context of a data-driven semiparametric nonlinear model. Similarly, Liu, (2022) developed an adaptive LASSO variable selection method for semiparametric spatial autoregressive panel models with random effects. The estimation is performed by maximising the concentrated profile likelihood function by means of a non-linear optimisation algorithm. Eventually, Chang et al., (2010) and Hsu et al., (2012) additionally reduced the model’s complexity by combining covariance tapering and PMLE for spatial and spatiotemporal settings, respectively.
@@
In addition to penalised likelihood methods, other estimation techniques have been developed for LASSO and its extensions in the geostatistical domain. For example, penalised least squares algorithms have been extended to the case of linear spatial models (Wang and Zhu,, 2009), spatial autoregressive models (Cai and Maiti,, 2020), regression models with spatially dependent data (Huang et al.,, 2010) and conditional autoregressive models (Gonella et al.,, 2022). Also, when considering additive spatial models with potential non-linear effects, weighted versions of penalised least squares can be applied (Nandy et al.,, 2017). Cai et al., (2019) proposed a generalised moments LASSO, which combines LASSO with GMM estimator, to perform variable selection for spatial error models with spatially autoregressive errors. Huang et al., (2010) propose using LASSO to simultaneously select relevant predictors, choose neighbourhoods, and estimate parameters for spatial regression with GIS layers to predict responses in unsampled sites. Safikhani et al., (2020) considered LASSO methods for generalised spatiotemporal autoregressive models. The estimators are obtained by a modified version of the penalised least squares that accommodates hierarchical group LASSO-type penalties.Chernozhukov et al., (2021) combine least squares LASSO and bootstrap procedures to get estimates and inference for systems of high-dimensional regression equations characterised by temporal and cross-sectional dependences in covariates and error processes. Cao et al., (2022) proposed a penalised estimation procedure for Gaussian Processes regressions where the likelihood and the first two derivatives are approximated by means of a scaled Vecchia approximation (Vecchia,, 2018). Eventually, several application-oriented papers combine classic LASSO approaches and geostatistical models in multi-step procedures (e.g., Fassò et al.,, 2022; Ye et al.,, 2011; Pejović et al.,, 2018).
@@
Penalised methods are also commonly applied in the context of functional data analysis, especially involving penalised splines (see Silverman and Ramsay,, 2002). These methods usually regularise the smoothness of the estimated functions by penalising the integrated second derivatives. In this way, many basis functions can be used, thus avoiding the typical overfit resulting from unpenalised estimation methods. Several authors have attempted to contribute by proposing LASSO-like penalised methods for selecting relevant functional (group) predictors (Pannu and Billor,, 2017) or to identify regions where the coefficient function is zero and to smoothly estimate non-zero values of the coefficient function (Centofanti et al.,, 2022). Basis expansions and low-rank representations (Wood,, 2017) are widely used tools in geostatistics for the spatiotemporal interpolation of environmental phenomena (see Hofierka et al.,, 2002; Xiao et al.,, 2016; Chang et al.,, 2010, for group-LASSO approaches in this context). For instance, in Maranzano et al., (2023), the authors make use of a PML estimator with adaptive LASSO penalty to select relevant functional covariates or their statistically relevant regions using the hidden dynamic geostatistical model. Eventually, Hsu et al., (2012) deal with semiparametric models for non-stationary spatiotemporal data in which a penalised least square with group-LASSO penalty is used to identify local spatial-temporal dependence features deviated from the main stationary structure.
3.2 Regularisation for geospatial interactions

### v2

@@ Abstract
This review article focuses on regularised estimation procedures applicable to geostatistical and spatial econometric models. These methods are particularly relevant in the case of big geospatial data for dimensionality reduction or model selection. To structure the review, we initially consider the most general case of multivariate spatiotemporal processes (i.e., $g>1$ dimensions of the spatial domain, a one-dimensional temporal domain, and $q\geq 1$ random variables). Then, the idea of regularised/penalised estimation procedures and different choices of shrinkage targets are discussed. Finally, guided by the elements of a mixed-effects model setup, which allows for a variety of spatiotemporal models, we show different regularisation procedures and how they can be used for the analysis of geo-referenced data, e.g. for selection of relevant regressors, dimensionality reduction of the covariance matrices, detection of conditionally independent locations, or the estimation of a full spatial interaction matrix.
Keywords: Spatial econometrics, geostatistics, variable selection, dimension reduction, big data.
@@ 1 Introduction
In the era of big geospatial data, the analysis of intricate spatial and spatiotemporal processes has become both increasingly vital and challenging. This is attributed to the growing volume and finer resolution of geo-referenced data, including remotely sensed, crowd-sourced or LiDAR data, as well as the increasing diversity of data types due to automated collection and processing, such as social network data and image data. These advancements enable the examination of more intricate relationships and allow for the detection of weaker dependencies or variations in the data. To statistically investigate such interactions, geostatistical and spatial econometric models offer a robust framework for comprehending the fundamental structures of these processes. Additionally, they provide insights into spatial dependencies, temporal dynamics, and the interactions among multiple random variables. An essential advantage of these statistical models lies in their interpretability, unlike deep learning models, which are often viewed as black-box models unless explicitly designed for interpretability. However, as the dimensionality of the spatial domain increases and datasets become larger and more complex, conventional modelling approaches frequently encounter challenges related to computational complexity and the appropriate selection of influential variables.
In this review paper, our focus is on regularised estimation procedures for spatiotemporal continuous data, mainly focusing on Gaussian geostatistical models and spatial autoregression models.The latter ones are often called spatial econometrics models, but we will refer to them as spatial autoregression models because the response variable is explicitly correlated with its adjacent observations in an autoregressive manner, and they are not only tied to applications in economics. Thus, these models are typically used for data defined on a discrete set of spatial locations, e.g., regular or irregular lattices, including polygons such as municipalities, counties or countries, making them attractive modelling approaches in econometrics (see, e.g., Billé and Arbia, 2019; Anselin, 2010; Arbia, 2016), ecology (see, e.g., Ver Hoef et al., 2018b ; Lichstein et al., 2002), or epidemiology (see, e.g., Lee, 2011; Gebreab et al., 2018). Instead, in geostatistics, spatial dependence is modelled using covariance functions depending on the distance between two observations [^2] Note that the distance is not necessarily measured as Euclidean distance, but different distance measures such as great-circle distances for modelling processes on a sphere or suitable network distances for processes on networks can be considered.. This makes the model attractive for processes on a continuous space and data which is irregularly observed across space, such as air pollution at ground-level measurement stations. Our focus will be mainly on stationary processes (across space and time) at a fixed set of spatial locations. Spatial point processes, such as Poisson processes, will not be the focus of this review paper (see González et al., 2016 for a review on spatial point processes).
@@
Regularised estimation procedures offer effective solutions for these challenges from multiple perspectives. For instance, they can be used for model selection and dimensionality reduction but also to reveal spatial dependence structures going beyond geographical proximity. This article will provide a comprehensive review of regularisation techniques, including shrinkage and penalisation methods, used for statistical modelling of geospatial data. Thereby, our focus will be on geostatistical models (Section 3) and spatial autoregression models (Section 4). Since the degree of regularisation is typically chosen based on predictive accuracy, we also review cross-validation techniques which are suitable under spatiotemporal dependence (Section 5). Finally, Section 6 concludes the article and discusses future research directions.
@@ 2 General framework
Assume that we are observing a continuous $q$-dimensional vector over space and time. The related $q$-variate dataset may be formalised as
$\{y_{t_{j}}(\boldsymbol{s_{i}})\in\mathds{R}^{q}:i=1,...,N,j=1,...,T\}$ (1)
where $\boldsymbol{s_{i}}$ are points in the spatial domain $D_{\boldsymbol{s}}$, which may be a Euclidean or non-Euclidean space. Common choices are the plain $\mathds{R}^{2}$, the Earth sphere $\mathds{S}^{2}$ or a discrete grid of points. The time domain, say $D_{t}\ni t_{i}$, is assumed discrete and, ignoring missing values, is made by equidistant time points. For simplicity, we use the set of integers, $D_{t}=\mathds{Z}$ and $t_{j}=j=1,...,T$. In the statistical framework, data $y_{t}(s_{i})$ are assumed to be generated by the spatiotemporal stochastic process $\{Y_{t}(\boldsymbol{s})\in\mathds{R}^{q}:\boldsymbol{s}\in D_{\boldsymbol{s}},t\in D_{t}\}$, (see Cressie and Wikle,, 2015). A continuous temporal domain is sometimes considered in geostatistics, see, e.g., Porcu and White, (2022), but we focus on the first case, which may be called the spatial time series approach.
The above-mentioned spatial and temporal domains can assume several forms, resulting in multiple data categories. In particular, following (Wikle et al.,, 2019, Section 2.1), we can distinguish among three main classes of spatiotemporal data. By geostatistical data, we mean phenomena which can be measured at continuous locations over a given spatial domain. Typical examples are air quality, weather and climate measurements, such as the PM10 concentrations reported in Figure 1. By areal or lattice data, we mean a phenomenon defined on a finite or countable subset in space over a specific time span. Examples of areal data are easily found in the socio-economic, political or medical-epidemiological realms, where observations are often reported by area (e.g., municipalities or regions). For example, in Figure 2, we show the evolution of per capita income for the European provinces (NUTS-3 classification) between 2011 and 2020. They provide a common playground for statisticians, econometricians, or applied social scientists, as the neighbourhood structure among units can provide relevant insights in many empirical contexts while preserving the interpretability of the phenomenon.
Figure 1: Illustrative example of real-world geostatistical spatiotemporal data. The plot shows the observed PM10 concentrations in 2018 at the 101 locations in Northern Italy considered by Fassò et al., (2023). For each panel, the observed average monthly concentrations are reported.
Figure 2: Illustrative example of spatiotemporal areal (irregular lattice) data. The plot shows the estimated yearly per capita gross domestic product (€ per capita) in the 1088 European Union provinces (NUTS-3 classification) from 2011 to 2020. Data source: Eurostat
There are different sources of correlation for such geo-referenced data:
@@
    Temporal proximity between two observations typically causes correlations. That is, the closer two observations are in the temporal domain $D_{t}$, the higher they are correlated, e.g., the current outside temperature will be similar to the temperature one hour earlier but less similar to the temperature two days ago.
    Also, more complex time-related structures are possible. For example, considering periodicity, the daily and seasonal cycles often require non-monotonic correlation structures.
@@
    Geographical proximity in $D_{\boldsymbol{s}}$ induces a similar dependence, known as spatial dependence. As claimed by Tobler’s first law of Geography, two observations are more similar if they are close to each other. For example, the current outside temperature in a city is similar to that in the neighbouring cities but less similar if you move away, especially in the North-South direction.
@@
There are two alternative, complementary approaches in spatial and spatiotemporal statistics to account for spatial dependence. Firstly, the dependence can be modelled using spatially dependent processes, where a suitable covariance function defines the entries of the covariance matrix. This approach is commonly known as geostatistics. Secondly, the dependent/outcome variable can be explicitly correlated with the nearby observations. Thereby, the local neighbourhood is defined by a suitable weight matrix $\mathbf{W}$, which is weighting all other observations, such that the product of $\mathbf{W}$ and the dependent variable is the weighted average of the adjacent observations. Below, we will provide an overview of regularised estimation procedures for both approaches, starting with geostatistical models in Section 3 and followed by spatial autoregression in Section 4.
@@ 3 Geostatistical models
A quite general frame for modelling data in (1) is given by the mixed-effects spatiotemporal model, which is given by
@@
where $\mu_{t}(\boldsymbol{s})$ is the fixed-effects component, $\omega_{t}(\boldsymbol{s})$ is the random-effects component, and $\varepsilon_{t}(\boldsymbol{s})$ is the stochastic error at time $t$ and location $\boldsymbol{s}$, which are typically assumed to follow a (Gaussian) white noise process. While the fixed-effects term typically models the mean behaviour and the influence of exogenous variables, the random-effects term accounts for any additional dependence in the data, including temporal and spatial dependence, heterogeneity, and cross-correlation between the variables. We will follow this distinction throughout the paper and discuss regularised estimation procedures for the parameters of $\mu_{t}(\boldsymbol{s})$ and $\omega_{t}(\boldsymbol{s})$. Penalised estimation procedures can be used to shrink certain model parameters towards a pre-specified target. The most often applied shrinkage target is zero. In this case, the parameters are shrunk towards zero, which means that a parameter is excluded from the model if its estimate is equal to the zero target. Thus, these methods are suitable for model selection (simultaneous parameter estimation and model selection). In particular, when applied to the fixed-effects part, they can automatically select relevant regressors. Compared to standard model selection procedures, such as step-wise selection based on cross-validation goodness-of-fit measures, regularised estimation procedures are typically computationally more efficient [^3] Note that only one or a few hyper-parameters, i.e., regularisation parameters, need to be selected via cross-validation, while for other complete or stepwise selection procedures, significantly more cross-validation iterations are needed..
Generally, there are two main sources of increasing computational complexity: model dimension and data dimension. The former is related to the fixed-effects component when considering variable selection. This is developed in Section 3.1. Moreover, the model dimension is related to the random-effects component when modelling spatial and temporal correlations, which are considered in Section 3.2.
Considering data dimension, the observation covariance matrix of the data set (1) is $qNT\times qNT$ dimensional, and its brute-force inversion has a cubic computational cost. Hence, specialised modelling and computing techniques are needed. The leading term in modern large spatial data sets is the number of spatial locations $N$. Fortunately, the correlation often decreases with an increasing distance between observations, allowing for a sparse covariance or precision matrix representation. Regularised estimation procedures with a shrinkage target of zero can be used to introduce zeros in the spatial covariance matrix or the precision matrix, indicating conditional independence between these observations. This will be the subject of Section 3.2. Similarly, if the number of variables $q$ increases, we observe a corresponding cubic increase characterised by the fact that the cross-correlation matrices for contemporaneous and colocated observations are usually dense. On the other side, the temporal dimension is often less critical because causal time series models may leverage on the fact that only past observations influence future observations, and the computational complexity may be reduced to be linear in $T$, see e.g. Wang et al., (2021), §2.4. These issues are also considered in Section 3.2.
3.1 Regularisation for the fixed-effects component
@@
Linear regression models, whether referring to spatiotemporal or unstructured data, can be estimated using various techniques, including Likelihood maximisation and least squares. However, in high-dimensional settings, traditional methods for regression models might not be directly applied (Nandy et al.,, 2017). In order to reduce the model’s complexity due to the fixed effect component $\mu_{t}(\boldsymbol{s})$, several methods for spatial and spatiotemporal settings have been proposed. Particularly, we refer to the task of selecting the most relevant predictors among a large set of candidates.
@@
The state-of-the-art statistical literature inherent to variable selection pivots around the Least Absolute Shrinkage and Selection Operator (LASSO) approach introduced by Tibshirani, (1996). LASSO is a penalised version of ordinary least squares that uses a LASSO-type $l_{1}$ penalisation to shrink irrelevant parameters to zero. Under mild regularity conditions, including uncorrelated observations, LASSO ensures consistent parameter estimation (see Section 2.4.2 of Bühlmann and Van De Geer,, 2011) and model selection consistency (Bickel et al.,, 2009; Belloni and Chernozhukov,, 2013), i.e., LASSO owns the oracle property. Among others, one of the most relevant features of LASSO is that it admits solutions to the minimisation task even when the number of parameters is greater than the number of available observations. Also, LASSO penalty (and extensions) can be used either considering least squares and penalised likelihood problems (Fan and Li,, 2001).
@@
It is worth noting that real-world geostatistical applications are prone to cross-correlated regressors due to their dependent (spatially and temporally) structure. Zhao and Yu, (2006) point out that when predictors are correlated, the classic LASSO algorithm does not provide selection-consistent estimates. Furthermore, when cross-correlation is detected, also group-LASSO estimators, which assume orthonormal data within each group, perform poorly in selecting the relevant predictors (Simon and Tibshirani,, 2012). A straightforward solution to this issue is provided by the adaptive LASSO penalty, which leads to selection-consistent estimators even in the presence of cross-correlated covariates (see Zou,, 2006) and Zou and Li, (2008).
@@
In addition to penalised likelihood methods, other estimation techniques have been developed for LASSO and its extensions in the geostatistical domain. For example, penalised least squares algorithms have been extended to the case of linear spatial models (Wang and Zhu,, 2009), spatial autoregressive models (Cai and Maiti,, 2020), regression models with spatially dependent data (Huang et al.,, 2010) and conditional autoregressive models (Gonella et al.,, 2022). Also, when considering additive spatial models with potential non-linear effects, weighted versions of penalised least squares can be applied (Nandy et al.,, 2017). Cai et al., (2019) proposed a generalised method-of-moments LASSO, which combines LASSO with GMM estimator, to perform variable selection for spatial error models with spatially autoregressive errors. Huang et al., (2010) propose using LASSO to simultaneously select relevant predictors, choose neighbourhoods, and estimate parameters for spatial regression with GIS layers to predict responses in unsampled sites. Safikhani et al., (2020) considered LASSO methods for generalised spatiotemporal autoregressive models. The estimators are obtained by a modified version of the penalised least squares that accommodates hierarchical group LASSO-type penalties.Chernozhukov et al., (2021) combine least squares LASSO and bootstrap procedures to get estimates and inference for systems of high-dimensional regression equations characterised by temporal and cross-sectional dependences in covariates and error processes. Cao et al., (2022) proposed a penalised estimation procedure for Gaussian Processes regressions where the likelihood and the first two derivatives are approximated by means of a scaled Vecchia approximation (Vecchia,, 1988). Eventually, several application-oriented papers combine classic LASSO approaches and geostatistical models in multi-step procedures (e.g., Fassò et al.,, 2022; Ye et al.,, 2011; Pejović et al.,, 2018).
@@
Penalised methods are also commonly applied in the context of functional data analysis, especially involving penalised splines (see Silverman and Ramsay,, 2002). These methods usually regularise the smoothness of the estimated functions by penalising the integrated second derivatives. In this way, many basis functions can be used, thus avoiding the typical overfit resulting from unpenalised estimation methods. Several authors have attempted to contribute by proposing LASSO-like penalised methods for selecting relevant functional (group) predictors (Pannu and Billor,, 2017) or to identify regions where the coefficient function is zero and to smoothly estimate non-zero values of the coefficient function (Centofanti et al.,, 2022). Basis expansions and low-rank representations (Wood,, 2017) are widely used tools in geostatistics for the spatiotemporal interpolation of environmental phenomena (see Hofierka et al.,, 2002; Xiao et al.,, 2016; Chang et al.,, 2010, for group-LASSO approaches in this context). For instance, Maranzano et al., (2023) use a PMLE with an adaptive LASSO penalty to select relevant functional covariates or their statistically relevant regions using the hidden dynamic geostatistical model. Eventually, Hsu et al., (2012) deal with semiparametric models for non-stationary spatiotemporal data in which a penalised least square with group-LASSO penalty is used to identify local spatial-temporal dependence features deviated from the main stationary structure.
3.2 Regularisation for the random-effects component